1. A ladder 26 feet long is leaning against the wall of a building. The ladder is pulled away from
the base of the wall at the rate of 3.5 feet per second.
a) How fast is the top moving down the wall when the base of the ladder is 10 feet from the wall?
b) Find the rate at which the angle between the ladder and the wall of the building is changing when the
base of the ladder is 10 feet from the wall.

2) Find the both absolute extrema for

f(x) = 145x2 – 4| on (-2,2].

please show work
Answer #1

A Here x and y are functions to t. TAC is the ladder of length 26 feet. AB i i the wall Them Pythagorean Theuns x2 + y2 = (26| Substitutionen der in substituting doze - 35 , x= 10,4+24 in ), 2x10x 3.5+ 2x 24 x dy zo dt 70 + 48 dy zo dt -70 48 ve signFrom faguse a cose non Coso dt We know that de - 3.5 and Y: 24, when x = lofest :dt .de - 21. 3.5 1 3.5 = 35 t = I dt 1 24 tof(2) - 0 + 45x3 - 4x = 0 x 145x2_4) 20 x = 0, 45x², 4 322 2 420, xz 2 x 2-2 315 are 3 butical points, - 315 Monotone InterwaWhen x E 10, 2 f (x) to . Tie flx) is Decreasing on 10, 2 ) 315) lie x20 is a local Maxima ( f (x) 70 befor x=0$ f(x)<o afte